OpenStudy (anonymous):

what is the slope of a line perpendicular to 2x - 3y = 12

OpenStudy (anonymous):



OpenStudy (anonymous):

its reciprocal is the perpendicular which is 3/2

OpenStudy (anonymous):

2x-3y=12 -3y=-2x+12 -(3y)/(-3)=-(2x)/(-3)+(12)/(-3) y=-(2x)/(-3)+(12)/(-3) y=(2x)/(3)-4 y=mx+b m=(2)/(3), b=-4

OpenStudy (anonymous):

prependicular, the slope is : -3/2